40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 9 12:24:46 2010, kiwamu, Update, SUS, strange ETMX suspension ETMX_spectr.png
    Reply  Thu Dec 9 22:04:40 2010, kiwamu, Update, SUS, Re:strange ETMX suspension 
Message ID: 4036     Entry time: Thu Dec 9 12:24:46 2010     Reply to this: 4038
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: strange ETMX suspension 

[Koji, Osamu and Kiwamu]

We found that the ETMX free swinging spectra showed a strange resonant frequencies.

We are going to inspect the suspension today.

 


(resonant frequencies)

In a ideal case the SOS (Small Optic Suspension) is supposed to have the following resonant frequecies.

(Although we didn't carefully identify which corresponds to which)

  f_POS ~ 0.98 Hz

 f_PITCH ~ 0.66 Hz

 f_YAW ~ 0.8 Hz

 f_SIDE ~ 0.99 Hz

However ETMX showed the following resonant frequencies.

  f_POS ~ 0.91 Hz

 f_PITCH ~ 0.7 Hz

 f_YAW ~ 0.93 Hz

 f_SIDE ~ 1.0 Hz

Especially f_YAW looks pretty high. Also the others are not at the right frequencies.

So we are suspicious that something wrong is happening on the ETMX suspension.

 

ETMX_spectr.png

ELOG V3.1.3-